5. A fair six sided die is rolled 10 times. Let X be the number of times the number 6 is rolled. Find P(X2)
Answer #1

Solution Assume 6 sreled due is fair zach haming grobability of 76. thus PCD = P(2) = P(s) = p(q) =P(A) = P(6) = 1/6. pex 52)